A man on the bank of the river pulls a boat on water with the rope that makes 30 degrees, with the bank. He exerts a force of 60lbs. What is the useful component of this force?why is it called useful?
1
Expert's answer
2020-10-25T18:29:06-0400

Solution

Diagram is shown as



Here is useful component

Is x- axis

"F_x=F\cos \theta
\\F_x=60\cos 30^°\\
F_x=60\times\frac{\sqrt{3}}{2}"

"F_x=51.96lb"

Change in Newton

"F_x=203.77N"